4x-3y = -10
5x-4=-13
Solve the system of equation

Answer by Alan3354(30924)

4x-3y = -10
5x-4=-13 --> 5x = -9
x = -9/5
----------
4*(-9/5) - 3y = -10
-3y = -10 + 36/5 = -2.8
y = 14/15